### v4

# EMNLP 2024 Paraphrase Types Elicit Prompt Engineering Capabilities

Abstract
Much of the success of modern language models depends on finding a suitable prompt to instruct the model. Until now, it has been largely unknown how variations in the linguistic expression of prompts affect these models. This study systematically and empirically evaluates which linguistic features influence models through paraphrase types, i.e., different linguistic changes at particular positions. We measure behavioral changes for five models across 120 tasks and six families of paraphrases (i.e., morphology, syntax, lexicon, lexico-syntax, discourse, and others). We also control for other prompt engineering factors (e.g., prompt length, lexical diversity, and proximity to training data). Our results show a potential for language models to improve tasks when their prompts are adapted in specific paraphrase types (e.g., 6.7% median gain in Mixtral 8x7B; 5.5% in LLaMA 3 8B; cf. Figure 1). In particular, changes in morphology and lexicon, i.e., the vocabulary used, showed promise in improving prompts. These findings contribute to developing more robust language models capable of handling variability in linguistic expression. Code: https://github.com/jpwahle/emnlp24-prompt-paraphrase.
1 Introduction
It’s not what you say it’s how you say it.
— Albert Mehrabian
Large language models (LLMs) already mimic human interaction by receiving instructions through natural language prompts and responding in natural language Radford et al. (2019); Ouyang et al. (2022); Touvron et al. (2023). The way prompts are designed has a marked impact on the value of a model’s output, and current LLMs require some degree of prompt engineering to be successful Liu et al. (2023b); Lu et al. (2023); Leidinger et al. (2023). A key step in prompt engineering is understanding how humans can express similar meanings differently, also known as paraphrasing. For example, consider the following paraphrases of a prompt, which share no words and vary greatly in length but convey the same message:
Avoid procrastination.
Stop postponing what you have to do.
Figure 1: The potential median task performance gain (blue) over the model’s baseline performance (orange) of five chat models across 120 tasks when their prompts were adjusted for specific paraphrase types (e.g., lexicon, syntax, morphology).
Figure 2: The main method of this paper. We paraphrase prompts of 120 tasks from 24 task families using 26 linguistic types of six categories (i.e., morphology, syntax, lexicon, lexico-syntax, discourse, and others) to analyze model inputs and outputs across different factors.
Humans understand and interpret the diversity in expressions, often without conscious effort. Arguably, LLMs should handle linguistic flexibility in a similar way to humans. Paraphrasing provides a window into the heart of prompt engineering; it gives us insight into what characteristics of the instructions language models value, what they understand, and where they lack capabilities. Ideally, LLMs should be robust to lexical, syntactic, morphological, inter alia, changes in the provided instruction — similar to how humans understand semantically identical texts written differently.
Through paraphrases, we can adjust prompts in these specific types to better understand which influence LLM’s ability to solve tasks. Emerging tracks in conferences and workshops on how to design prompts underline a substantial interest in understanding which nuances of prompts affect models Graux et al. (2024).
To determine the influence of linguistic types on prompts, we need a shared understanding of the variations between similar prompts. Vila et al. (2014) introduced paraphrase types which define linguistic changes at individual text positions. Wahle et al. (2023) adopted paraphrase types to train models that generate paraphrases with these capabilities. Paraphrase types can identify linguistic types, such as whether the lexicon has changed (by replacing words) or the syntax has changed (by altering grammatical structure).
So far, there has been little work to systematically evaluate how linguistic variations of prompts affect current language models. We fill this gap by evaluating hundreds of thousands of prompt variations across 120 tasks and five models (3.24m total prompt-example combinations). Our procedure is outlined in Figure 2. We measure how sensitive the models are to different prompt changes and what we need to improve their robustness and effectiveness in performing different tasks. Our work addresses the following key questions:
  1. 1.
    How sensitive are the models to specific paraphrase perturbations of the prompt?
  2. 2.
    Which linguistic changes in the prompt affect models the most?
  3. 3.
    What other factors play a role, such as the length and lexical diversity of the prompt or proximity to the training data?
  4. 4.
    How does the above evolve across models and tasks?
Our results show that adapting prompts to specific types (e.g., morphology) can yield significant gains in many models (cf. Figure 1) across different downstream tasks, such as summarization, text completion, or dialogue generation. We also control for various confounding factors, such as the length of prompts, lexical diversity, and proximity to training data, and show that performance gains can be observed independent of these additional factors. We recommend which paraphrase types to consider when adapting prompts for a particular task (e.g., polarity substitutions for sentiment analysis).
2 Related Work
Prompt-based learning has become a new trend, where pre-trained language models perform prediction using a template description of the intended task, and the model derives the necessary information without the need for gradient updates Liu et al. (2023a). Prompt-based learning has advantages over traditional supervised learning, as the same pre-trained LLM can perform different unseen tasks without fine-tuning. Early prompt tuning methods focused on integrating trainable continuous prompt embeddings to perform various NLU tasks Gu et al. (2022); Liu et al. (2021, 2023b). Discrete prompts representing actual tokens have become more popular than continuous embeddings in prompt engineering, arguably because they are more accessible and linguistically interpretable.
Discrete methods used cloze-style phrases and differentiable prompts to improve few-shot learning capabilities Schick and Schütze (2021); Zhang et al. (2022). Shin et al. (2020) proposed AutoPrompt, which uses a masked language model to find variations of prompts and has been used for various downstream tasks such as plagiarism detection Wahle et al. (2022). Zhou et al. (2022) introduced Automatic Prompt Engineer, which searches over a pool of prompt candidates proposed by a second LLM. Other work focused on generating prompts for knowledge extraction from LLMs Jiang et al. (2020) or investigated instruction induction and model tuning with minimal human intervention Honovich et al. (2022b, a). Recent work uses variations of Self-Instruct to improve model prompting Wang et al. (2022a).
Reynolds and McDonell (2021); Hu and Levy (2023) evaluate the role of prompts in model control and the limitations of metalinguistic prompts for assessing linguistic knowledge in LLMs. Studies by Leidinger et al. (2023) and Lu et al. (2023) examined the influence of linguistic features of prompts on LLM performance with hand-selected prompt variations (e.g., tense, modality). Mizrahi et al. (2024) demonstrated through empirical analysis that LLM performance can vary based on how task instructions are phrased. Sorensen et al. (2022); Yang et al. (2023) explored the optimization of prompt selection using unified scoring frameworks and unsupervised techniques based on mutual information.
Our work contributes to previous work in several key aspects. While previous studies have explored hand-crafted linguistic features on LLM performance Leidinger et al. (2023); Lu et al. (2023), we take a systematic approach by decomposing paraphrases into a set of six families of changes (i.e., syntax, lexicon, lexico-syntax, morphological, semantic, and others). Although other works have proposed strategies for prompt selection Yang et al. (2023); Sorensen et al. (2022) and methods for enhancing prompt tuning and generation Liu et al. (2021); Shin et al. (2020); Zhou et al. (2022), our research quantitatively measures the effects of linguistic paraphrase types on model responses in a bottom-up method to apply successful types on different tasks. Other work focuses on domain-specific problems with few tasks. Our methodology spans five models, 120 tasks, and 24 task families. The main gap we address is the lack of empirical evaluation of how different linguistic manipulations of prompts affect LLM behavior and performance on a large scale. Our work controls for other confounding factors, such as prompt complexity, training data proximity, and lexical diversity.
3 Data & Models
Central to this study is a dataset of tasks, prompts, and variations of these prompts. We use the Super-NaturalInstructions dataset Wang et al. (2022b), which contains more than 1 600 tasks with their respective prompts. We describe in detail how we construct different variations of these prompts in the experiments (Section 4). We sample 120 different tasks (each task has its own dataset) from 24 different task families (e.g., question answering, sentiment analysis) with the following conditions: the task family must have at least 10 tasks, the primary language is English, and each task has 200 examples to provide sufficient statistical power. We sample 5 tasks per domain, leaving us with 24 task families x 5 tasks per domain = 120 total datasets. See Appendix A for the task details.
We choose the five best non-proprietary models according to the LMSYS chatbot arena leaderboard in descending order of performance as of 1 May 2024111https://chat.lmsys.org/?leaderboard: LLaMA 3 Instruct (70B), Command R+ (104B), Mixtral 8x7B Instruct (47B), LLaMA 3 Instruct (8B), and Gemma Instruct (7B). Unless otherwise noted, we set the temperature to 0.2, the probability mass to sample words (top p) to 0.9, the penalty for repeatedly sampling the same sequence of tokens to 0.1, and the maximum number of tokens generated to the average of human references for that task. This number varies for different tasks but averages between 1 and 3154 tokens (see Appendix A for details). We use 40 NVIDIA A100 GPUs (40GB) with 16-bit precision (8 A100s for each model). Our experiments required a total computational budget of approximately 2880 A100 GPU hours, resulting in 311kg of CO2 equivalent [^2] https://mlco2.github.io/impact/#compute.
4 Experiments
Q1. How do different linguistic changes in the prompt affect the performance of a language model? Which changes affect a model the most?
Ans. We start with a seed prompt designed for a specific task, e.g., “In this task, you are given Twitter posts. Your task is to label the post’s emotion as sadness, joy, disgust, anger, fear, or surprise.” We paraphrase each prompt with one of 26 linguistic changes (e.g., syntax, lexicon) using the gpt-3.5-turbo-16k model, which we fine-tuned in the same way as Wahle et al. (2023) using the ETPC dataset Kovatchev et al. (2018). See Table 9 and Figures 8 and 10 in Appendix A for an extensive list of types, full prompt template and examples respectively. For the original prompt and its 26 paraphrases, we test each model’s few-shot settings. Depending on the task, the number of few-shot examples ranges from 3 to 6 (sum of positive and negative examples).
Figure 3: The average downstream task performance gain or loss from applying specific paraphrase types to the prompt for all 120 tasks and five models.
Results. As Figure 1 already revealed, we can observe marked changes when adjusting prompts. Command R+ experiences up to 6.3% median performance gain, and Gemma 7B up to 13.4%. We decompose these results by the different types in Figure 3. Lexicon changes (+1.26%), closely followed by syntax changes (+1.19%), account for the largest median performance gain. The potential loss is comparable across changes. Overall, paraphrases had more upside than downside potential (+1.16% vs. -0.87% median change).
Discussion. Across all tested models, there seems to be marked upside potential when we adapt prompts in specific linguistic types. Morphology changes include changing modal verbs, which helps LLMs to follow instructions more clearly, e.g., “In this task, one should must detect the sentiment of the sentence.” Lexicon changes have shown success across our experiments; in particular, we found examples in which more specific vocabulary benefits a task prompt, e.g., instead of “Determine how people feel about this text.” a more precise version yielded better results: “Determine whether the sentiment expressed is positive, negative, or neutral”,
Q2. Are there prompt changes consistently improving performance on a particular set of tasks? Do these changes have a different impact on tasks in different domains?
Ans. We decompose the results of Q1 into individual task families, i.e., a set of related problems, such as sentiment analysis or question answering. A task is a specific set of data and instructions for solving a problem in a task domain, e.g., classifying emotions on X (formerly known as Twitter).
Figure 4: The avg. gain or loss in performance for all 120 tasks in 24 different task families across all five models.
Results. Figure 4 shows that title generation (+6.01%), text completion (+5.86%), and question answering (+5.60%) gain the most performance while having a low potential for a negative impact. Commonsense classification (-4.86%), sentiment analysis (-4.83%), and word semantics (-4.82%) have the highest loss potential.
Specific perturbations of the prompt affect the performance of models on tasks from varying domains differently. Morphology changes in the prompt have the largest gains in wrong candidate generation (+26.0%), question generation (+21.5%), and textual entailment (+17.5%). Lexical changes show consistent gains in summarization (10.8%), wrong candidate generation (+8.9%), and title generation (+7.1%). These tasks rely on semantic precision and vocabulary richness to interpret or generate nuanced responses. Positive effects with discourse changes can be observed in tasks that require an understanding of longer contexts or multi-sentence structures, such as summarization (+7.8%). Detailed results for each model are available in Tables 10, 12, 13, 11, 15 and 14 in Appendix C.
Discussion. Some tasks are more sensitive to instructions than others. Tasks such as question generation and text completion may benefit from syntax changes because they can improve grammatical accuracy while reducing the flexibility of more creative tasks such as dialog generation. In particular, sentiment classification seems to benefit from polarity substitutions and is sensitive to negation, consistent with related work Wiegand et al. (2010). Lexical changes improve performance in vocabulary-intensive tasks such as named entity recognition and text entailment. However, they may hinder performance in more general contexts such as question rewriting, which is often based on data from Wikipedia and knowledge bases (see Table 5 in Appendix A). In discourse, better context management, as well as indirect and direct style, improves clarity and coherence in tasks like summarization and long-form answer generation.
Sometimes, even the same changes can produce different results for different models. For example, larger models such as Command R+ benefit from morphology changes for summarization (+23.6%), while the same changes result in a loss of performance in the two small models LLaMA 3 8B (-16.4%) and Gemma 7B (-16.1%).
Q3. How sensitive are different models and architectures to changes in the prompt? Are there differences between models of different sizes? Can a smaller model achieve better results than a larger model by paraphrasing its prompts?
Ans. Related work has shown that model size and training scale play a central role in language model success, also known as “scaling laws” Kaplan et al. (2020); Rae et al. (2021); Wei et al. (2022). What has not been shown so far is whether and how much lexical changes influence different model architectures and sizes when adapting prompts. Specifically, we investigate how to bring the performance of a weaker model up to the level of a larger model by tweaking its prompts.
Results. In Figure 1, note how the smaller 8B version of LLaMA 3 can gain up to 5.5% median task performance, while the larger 70B model can gain much less with up to 2.8%. LLaMA 3 8B has a lower baseline performance of 0.58, while the 70B model’s baseline is at 0.65 (see Figure 12 in Appendix C for more details).
LLaMA 3 models are less sensitive to changes than other models (0.05 avg. std. in LLaMA models vs. 0.10 avg. std. over other models). Smaller models are more sensitive and have the highest possible gain potential when adjusting prompts (e.g., Gemma 7B: +10.2% gain through lexicon; LLaMA 8B +8.2% gain through morphology). The most sensitive model in our experiments is Command R+ (avg. std. 0.16).
Although Command R+ achieves an average score of 0.61, which is lower than LLaMA 3 70B’s with 0.65, it can outperform LLaMA 3 70B by 0.08 when its prompts are tuned. The same applies to Mixtral 8x7B (0.55) and LLaMA 3 8B (0.58). Mixtral’s prompts, when tuned, can score 0.06 higher than the previously better LLaMA model.
If we always choose the best (paraphrased) prompts, we can achieve even higher performance gains; for example, LLaMA 3 8B could gain 21.1%, making the model markedly better than its 70B counterpart by tweaking the prompts. We want to note that it is difficult to always find the best prompt. Therefore, we report median performance gains in the main body of the paper and report results for selecting the best prompt in Figure 9 in Appendix C.
Discussion. Models have different architectures and training processes, especially training data, which play a fundamental role in their behavior. It is encouraging to see that there is still upside potential to improve task results without relying on more computational resources for training. However, this raises the question of why such different behaviors can be achieved by changing the instructions. An interesting parallel is how humans can also sometimes produce different results depending on the instructions given. Our results suggest that smaller models can perform similarly to larger ones and are more sensitive to paraphrase changes.
Q4. Do paraphrased prompts that increase a model’s task performance also show greater similarity to the model’s training data?
Ans. We know that a confounding factor in prompt engineering is that prompts that are closer to a model’s training data improve their likelihood of answering more confidently Zhao et al. (2021). We use the FineWeb corpus with 350 billion tokens ($\approx$388GB) to search for examples close to paraphrased prompts and compute their similarity.
We build a BM25 index ($\approx$610GB), and for each prompt and its 26 variations, we query this index to find the closest examples. We measure the difference in similarity of the original prompt to the training data versus our paraphrased prompts to the training data by computing the following measure:
$\Delta_{train}=RL(P,T_{P})-RL(O,T_{O})$ (1)
where $RL$ is the ROUGE-L score, $O$ is the original prompt and $T_{O}$ is the nearest training example for $O$ (as measured by the BM25 score), $P$ is the paraphrased prompt, and $T_{P}$ is the closest training example for $P$.
Figure 5: The distribution of how much closer the paraphrased prompt is to the closest training example in FineWeb 350BT over the original prompt (x-axis) and the distribution of task performance (y-axis). Red colors mean high mass and blue colors mean low mass between $\Delta_{train}$ and task performance.
Results. The x-axis of Figure 5 shows $\Delta_{train}$, which represents this difference in similarity to the training data between the paraphrased and original prompts. The y-axis then shows downstream task performance. The results show few very close training examples for our tasks, i.e., the ROUGE-L between prompt and closest FineWeb example is typically below 0.5. Most successful paraphrased prompts with a downstream task performance greater than 0.8 do not show higher similarity to training than the original prompt. Sometimes, higher scores can be achieved when paraphrases are closer to the training (top right of the figure).
Discussion. Consistent with related work Zhao et al. (2021), we show that prompts closer to training examples sometimes have higher accuracy than those without close training examples. However, this does not generally hold across our large set of tasks. We found no evidence that paraphrases were better because they had closer training examples.
Figure 6: The percentage in task performance gain and lexical diversity as measured by four metrics in four tasks. The light blue quadrants show areas of high lexical diversity and strong task performance gains.
Q5. How do different prompt perturbations affect the lexical richness of language model responses in generative tasks?
Ans. We have previously measured downstream performance as a key indicator of prompt quality. In classification tasks, this is directly measurable by the binary downstream metric (i.e., “yes” or “no”). In generative tasks, it is not easy to say whether a generative answer is correct using a human reference Clark et al. (2021); in particular, an output may be correct even without having any words in common with a reference (recall our example from the introduction). Therefore, we can evaluate complementary factors about responses (e.g., whether the response is creative or uses lexically diverse language). A key factor in complex responses in natural text is lexical richness Van Hout and Vermeer (2007); Jarvis (2013); Kyle (2019).
We measure lexical richness with four metrics that focus on the diversity of tokens and text segments Guiraud (1960); Maas (1972); Covington (2007); McCarthy and Jarvis (2010): Root Type-Token Ratio (RTTR), Maas, Moving Average Type-Token Ratio (MATTR), and Measure of Lexical Diversity (MTLD) — more details in Appendix B. We select four task families with long responses to measure lexical richness: summarization, question generation, question rewriting, and dialogue generation (20 tasks in total).
Results. Summarization tasks have a particularly high diversity in responses as measured by Maas, RTTR, and MTLD (Figure 6 (a), (c), (d)). Overall, lexicon and lexico-syntactic changes often lead to higher performance at the expense of lexical richness. Models produce responses with lower lexical diversity in question rewriting tasks as measured by MAAS (Figure 6 (a)). Syntax and other changes lead to both lower lexical diversity and task performance gains compared to others measured by RTTR (Figure 6 (c)). Question rewriting achieves the highest lexical diversity as measured by MAAS, specifically with lexicon and lexico-syntactic changes (Figure 6 (a)), and question generation and question rewriting have overall lower diversity and performance gains, specifically with discourse and other changes (Figure 6 (c)).
Discussion. Unexpectedly, summarization and question rewriting, which are arguably less open-ended and based more on extracting or paraphrasing information, lead to high lexical diversity. Purely generative and more open-ended tasks such as question generation and dialogue generation yield lower lexical diversity scores. Larger changes in the prompt markedly affect the lexical diversity of model responses. This is not surprising, as asking people to perform tasks in very different ways often leads to more diverse responses. What is surprising, however, is that marked performance gains can still be observed. Since these changes lead to substantial variance in the language model output, our results suggest that the complexity of the prompt (e.g., its length, word positions, or lexical shifts) can also be a confounding factor to task performance.
Q6. What role does prompt complexity play in task performance outcomes? To what extent do variations in length, word position, and lexical variation correlate with downstream task performance?
Ans. We use three markers to describe the complexity of a prompt paraphrase relative to the original instruction: the deviation in absolute number of tokens, word position deviation, and lexical deviation Liu and Soh (2022). We use Pearson correlation to measure a quantitative relationship between these three markers of prompt complexity and task performance. More details about the metrics and correlations can be found in Appendix B.
Figure 7: The performance difference between original and paraphrased prompts for temperatures from zero to one.
Results. Across tasks, there is no significant correlation between changes in prompt length, changes in word position, and lexical deviation, as all correlations are close to zero (see Table 2 in the Appendix C). We found no evidence that making prompts more or less complex is associated with higher or lower task performance.
Discussion. Note that when paraphrasing prompts in certain lexical types, it is not new complexity added to the prompt that leads to different results (both gains and losses in task performance). This suggests that the actual linguistic perturbations make the prompt easier for the model to understand, leading to higher scores. However, there may be other factors that can also play a role.
Q7. How much randomness is involved in these changes? Can the models’ temperature lead to particularly good or bad task results that partially explain these gains or losses?
Ans. When prompting models to perform tasks, we choose 11 temperature variations for LLaMA 3 70B from zero to one with steps of 0.1. For title generation and summarization, we run the original prompt and its paraphrased version using morphology and lexicon changes with these different temperature variations. We do this 11 times, with each run having a different temperature for the same paraphrased prompt, and then compute the difference between task performance of the original and paraphrased versions as:
$\Delta_{temp}=RL(LM(O_{i}))-RL(LM(P_{j}))$ (2)
where $RL$ is the ROUGE-L score, $O_{i}$ is the original prompt with temperature $i$, $P$ is the paraphrased prompt with temperature $j$, and $LM$ is the language model output.
Results. Figure 7 shows contour plots of different temperature choices. The coloring shows $\Delta_{temp}$ for different temperatures for the paraphrased (x-axis) and original (y-axis) prompts. Note how the distributions within the summarization have similar hills and valleys (see Figure 7 (c) and (d)). For title generation, with generally higher performance loss, the highest gains are achieved for low to medium temperatures (up to 0.6 for the paraphrased prompt). For summarization, with overall higher scores, low temperature yields the highest performance difference, close to 0 for the paraphrased prompts and between 0.2 and 0.4 for the original.
Discussion. While we cannot fully exclude that randomness might improve results (i.e., there are some hills even in the high temperature ranges), most of the gains stem from low temperatures, suggesting that significant performance gains are due to paraphrase types. It may also be worth mentioning that additional randomness can come from the models trained by Wahle et al. (2023), as they can only paraphrase with some degree of accuracy.
5 Conclusion
This study evaluated five language models across 120 tasks (including sentiment, question answering, commonsense reasoning, summarization, etc.) and showed that paraphrasing prompts can improve the performance of language models. We showed that language models have a marked upside potential to improve task results when their prompts are adapted in specific linguistic types (e.g., polarity substitutions for sentiment analysis). Other work can use our findings about which tasks benefit from which paraphrase types to design new prompts. We also controlled for prompt complexity, temperature, and proximity to training data.
Current model performance represents a lower performance bound for tasks as we showed that semantically identical instructions hold marked upside gain. While it is not entirely clear why language models are often sensitive to changes in instruction, we have systematically tested different lexical features and found that some have a larger positive impact (e.g., morphology) than others (e.g., syntax), depending on the tasks. Since humans understand tasks presented in different ways and are robust to small (or even complex) changes in instruction, language models should have a similarly robust interface to communication in the future. We have contributed to this development of robust language interfaces by showing how specific types can benefit or harm models over a large set of tasks and prompts. One can also use our approach to create prompts the model does not understand to augment training data and increase its robustness.
Limitations
Our results show that the same paraphrase changes can potentially improve results for one model but harm another. The same is true for different changes in the same task. In a task like sentiment, sometimes the same polarity substitutions lead to improvements, and sometimes they do not. Variance across examples and models does play a role here but another possible reason is that models trained to generate paraphrase types only have a certain accuracy, leading to models sometimes confusing one type for another or generating an incorrect type as recent work shows Meier et al. (2024). However, at a large aggregate level (3.24 million prompt-example combinations), our results provide important trends about prompts despite these margins of error. Better models for the controlled paraphrase generation will provide more accurate results in the future.
Our selection of paraphrase types is not a complete set of all flexibility in linguistic expression. There may be other variations, especially extremes, that we have not considered. However, the set of paraphrases across morphology, lexicon, lexico-syntax, syntax, and discourse covers many of the most common paraphrases encountered in texts. We also find that there is no single paraphrase type that improves a model’s accuracy across tasks consistently. This is somewhat to be expected, as different tasks may benefit from different adaptations even though the overall domain is the same. The same is true for tasks in different domains; some modifications are successful in one domain but not in another. For example, polarity and negation play a larger role in sentiment, while lexical changes affect vocabulary intensive tasks (e.g., named entity recognition) or tasks that require specificity in vocabulary (e.g., example). Finding the most successful type of change in a given setting is non-trivial, and more research needs to be done on successfully perturbed prompts for new and unseen tasks.
Acknowledgements
This work was partially supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. Many thanks to Andreas Stephan and Lars Käsberg for their thoughtful discussions.
Appendix A Details on the Prompts & Tasks
Prompts. We use the template shown in Figure 8 to construct prompts for each of the 120 tasks. In Figure 10, we show an example for the numerical commonsense reasoning task NumerSense in which the model has to predict a blank token ‘_’ Lin et al. (2020). To paraphrase prompts, we use one of 26 paraphrase types in six groups shown in Table 9. Five examples of paraphrases applied to prompts can be seen in Table 1. In the first case, the added specification made the instruction more precise; in the second case, the model removed seemingly unnecessary context. The third case involved replacing words or phrases with synonyms or near-synonyms that have the same meaning in the given context, while the fourth case involved restructuring the sentence or changing the way the information is presented while maintaining the same overall meaning. Finally, the fifth case replaced a single word with a phrase (or vice versa) that conveys the same meaning. Further paraphrase type definitions and perturbation examples can be found in Vila et al. (2014); Meier et al. (2024).
Tasks. We provide an overview of all 120 tasks together with the paraphrase type that, when applied to prompts in that task domain, had the most positive influence together with the potential downstream task performance gain over the original prompt in Tables 4, 5, 6 and 7. Further, we want to note an observation of our experiments showing that tasks containing content judged as unethical (such as toxic language detection on Twitter) led to many denied responses by the chat or instruction-finetuned models. Likely because of their post-training, models answered that they could not respond to these questions, although we did not ask them to produce new harmful content but to classify existing content. We removed these tasks from our final analysis. This observation raises questions about how many powerful models can be used in the future to classify toxic content when their guardrails prevent them from answering.
Task Prompt Template <system>
     You are a helpful assistant in solving various tasks. You should only output one answer to the task, nothing more,
no explanations, and nothing around. Just read the instruction carefully, understand the positive and negative examples provided, and generate one single answer in the same way as the example’s output. </system>
<user>
     Instruction: {instruction}
Positive examples: {positive examples} Negative examples: {negative examples} Input: {example} Output:
</user>
Figure 8: The prompt template used to perform each of the 120 tasks. For Mixtral 8x7B Instruct and Gemma 7B Instruct, we prepend the system prompt to the user prompt because the models do not natively support system prompts.
Figure 9: The maximum performance gain models can experience if their prompts are adjusted in optimal ways as explained in Table 8.
Prompt Example for Numerical Commonsense Reasoning <system>
     You are a helpful assistant in solving various tasks. You should only output one answer to the task,
nothing more, no explanations, and nothing around. Just read the instruction carefully, understand the positive and negative examples provided, and generate one single answer in the same way as the example’s output. </system>
<user>
     Instruction: Find the most appropriate number to replace the blank (indicated with _) and express it in words.
Positive examples:
          Input: A lion has _ legs.
Output: four Input: Numbers less than _ are called negative. Output: zero Input: There are _ hours in a day. Output: twenty four Negative examples:
          Input: A dog has _ legs.
Output: 4   # Not expressed in words but numbers. Input: Numbers less than _ are called negative. Output: one   # Logically wrong. Input: Some plant varieties can grow up to _ feet tall. Output:
</user>
Figure 10: A prompt example for the first instance in the NumerSense dataset Lin et al. (2020).
Instruction Type Instruction Text ROUGE-L
Example 1: Addition/Deletion
Original Your task is to generate a headline (title) for this article. 0.60
Paraphrased Your task is to generate a concise headline (title) for this article. 0.78
Example 2: Ellipsis
Original Your task is to generate a headline (title) for this article. 0.60
Paraphrased Generate a headline (title) for this article. 0.67
Example 3: Same Polarity Substitution (contextual)
Original Summarize the main points of the given text in 3-4 sentences. 0.55
Paraphrased Condense the key ideas of the provided passage into 3-4 sentences. 0.72
Example 4: Syntax/discourse structure changes
Original Explain the concept of gravity as if you were teaching a 10-year-old child. 0.63
Paraphrased Imagine you’re teaching a 10-year-old child. Explain the concept of gravity. 0.71
Example 5: Synthetic/analytic substitution
Original Translate the following English text into French. 0.58
Paraphrased Convert the given English passage to its French equivalent. 0.70
Table 1: Comparison of original and paraphrased instructions for five types with downstream ROUGE-L scores.
Appendix B Details on the Metrics
We provide the mathematical equations and interpretations of metrics we used to compute various aspects of this study, such as lexical diversity, prompt complexity, potential gains, etc. in the following subsections.
B.1 Prompt Complexity Metrics
Absolute Token Deviation. An intuitive heuristic that measures the total number of tokens added, removed, or changed between two text pairs. This metric reflects the overall extent of alteration in content and can be computed by:
$tok(s1,s2)=|N_{s1}-N_{s2}|$ (3)
where $N_{s}$ represents the number of tokens in sentence $s$. This metric captures the absolute difference in the length of the texts, offering a simple yet effective way to quantify the overall size of changes.
Word Position Deviation. This metric measures structural alterations by calculating the average shift in the positions of common words between two paraphrased sentences. The equation uses the mean of the maximum relative position shifts of all common words between two sentences $s1$ and $s2$, represented as
$pos(s1,s2)=$ (4)
$\frac{1}{N_{\mathcal{C}}}\sum$ $max\{\delta_{s1,s2}(W),\delta_{s2,s1}(W)\}$ (5)
where $\delta_{s1,s2}$ is the relative position shift of a word $W$ with respect to sentence $s1$ in paraphrase pair $(s1,s2)$, and $NC$ is the count of common words.
Lexical Deviation. This measure quantifies the difference in vocabulary used between two sentences. It is defined as the complement of the ratio of the number of common words to the total number of unique words in both sentences, given by the equation
$lex(s1,s2)=1-\frac{N_{\mathcal{C}}}{N_{\mathcal{A}}}$ (6)
where $N_{\mathcal{A}}$ is the count of all unique words that occur in either occurs in either or both sentences and $N_{\mathcal{C}}$ is the set of common words that occur in both sentences.
B.2 Prompt Complexity Correlations
We calculate the Pearson correlation between each of the above prompt complexity metrics and the downstream performance in the following way.
$P_{0}$ represents the original prompt and $P_{1},P_{2},\ldots,P_{26}$ denote the paraphrased versions. The performance score for each paraphrased prompt $P_{i}$ is given by $S_{i}$, where $S_{i}$ is ROUGE-L in this study. For each paraphrase $P_{i}$, the marker of change relative to $P_{0}$ is denoted by $T_{i}$ ($T_{1}$ being deviation in absolute number of tokens, $T_{2}$ word position deviation, and $T_{3}$ lexical deviation). We calculate Pearson correlation $r$ between $T_{i}$ and task performance $S_{j}$:
$Pea$ $rson(i,j)=$ (7)
$\frac{\sum_{i=1}^{n}(T_{i}-\bar{T})(S_{i}-\bar{S})}{\sqrt{\sum_{i=1}^{n}(T_{i}-\bar{T})^{2}}\sqrt{\sum_{i=1}^{n}(S_{i}-\bar{S})^{2}}}$ (8)
where $n$ represents the number of paired scores available (26 paraphrase-original pairs), $\bar{T}$ is the mean value of all $n$ markers of change, and $\bar{S}$ is the mean value of all $n$ performance scores.
B.3 Lexical Diversity Metrics
We use four main metrics to assess lexical richness for LLM generations: Root Type-Token Ratio (RTTR), Maas, Moving Average Type-Token Ratio (MATTR), and Measure of Lexical Diversity (MTLD). RTTR and Maas metrics consider the entire set of texts for a given category, focusing on the number of distinct words relative to the total number of words, with modifications to account for text length.
RTTR. Root Type-Token Ratio measures the proportion of unique words in a text, normalized by text length, indicating the lexical variety of the text Guiraud (1960). Computing the quotient of distinct words and the total number of tokens, the metric is defined as:
$\text{RTTR}=\frac{T}{\sqrt{N}}$
where $T$ represents the number of unique types (distinct words), and $N$ is the total number of tokens (words).
Maas. This metric measures the lexical richness by considering the relationship between the number of distinct words and the total word count, using a logarithmic transformation to reduce the impact of text length Maas (1972). The measure is calculated using the formula:
$\log(V)=\log(N)+\alpha\log(N)$
where $V$ is the number of distinct words, $N$ is the total number of words, and $\alpha$ is a parameter calculated as follows:
$\alpha=\frac{\log(V)-\log(N)}{\log(N)}$
MATTR. Moving Average Type-Token Ratio measures the stability of lexical diversity across different text segments, providing an average diversity score that accounts for variations within the text Covington (2007). Taking the average Type-Token Ratio (TTR) over a sliding window of size $w$, which in our experiments is 25 tokens, is calculated as:
$\text{TTR}_{i}=\frac{T_{i}}{N_{i}}$
where $T_{i}$ is the number of unique types within the window $i$, and $N_{i}$ is the number of tokens within the window $i$. The MATTR is then the average of these TTR values over all windows:
$\text{MATTR}=\frac{1}{m}\sum_{i=1}^{m}\text{TTR}_{i}$
where $m$ is the number of windows.
MTLD. Measure of Lexical Diversity determines the text’s lexical diversity by determining how many words are needed before the diversity drops below a set threshold, thereby capturing the consistency of word usage variety throughout the text McCarthy and Jarvis (2010). MTLD measures the text length until the Type-Token Ratio (TTR) reaches a threshold (0.72 in our experiments), then starting a new segment and continuing this process through the entire text measures MTLD. Mathematically:
MTLD=Total tokensNumber of $\text{MTLD}=\frac{\text{Total tokens}}{\text{Number of segments}}$
where a segment is defined as the number of tokens until the TTR reaches the threshold.
Higher RTTR (ranging from 0 to the square root of the total number of tokens), MATTR (0 to 1), and MTLD (typically 0 to infinity) values indicate greater lexical diversity, while lower values suggest less variety in word usage. The Maas metric, which typically ranges close to 0, operates inversely: lower values indicate higher lexical richness, while higher values suggest lower diversity.
B.4 Downstream Task Metrics
Because we evaluate on the benchmark suite of Wang et al. (2022b), we use the corresponding evaluation metric ROUGE-L for each task. Their study shows that ROUGE-L correlates well with human judgments for these tasks and yields comparable scores for all generative tasks. Further, this choice requires no postprocessing or judgment by another model, which could have its own biases.
B.5 Potential Downstream Gain
Figure 1 has shown the potential median gain that can be reached by paraphrasing prompts in specific types. In Table 8, we show how the median is calculated per example. For each example in a task dataset, we take the downstream task scores of paraphrased prompts that are higher than the original prompt and take the median. Then, for all examples in the dataset, we average that median.
Further, one can also compute the maximum performance that can be reached if the best paraphrase type would be chosen to paraphrase the prompt. As previous experiments have shown, it is challenging to devise a robust method to select the best paraphrase of a prompt. More research on the Monte-Carlo approach could bring us closer to this maximum. The median is a fair way to represent a gain that users could expect by selecting one of the multiple paraphrases of prompts that are better than the original. Compared to the mean, the median is not influenced by outliers that are unlikely to select (e.g., the two 1.0 scores in Table 8 are unlikely to reach; the most likely expectation is 0.67, which is also the median).
Task Family ($\downarrow$) Tok. Pos. Lex.
Answerability Class. -0.06 -0.06 -0.06
Commonsense Class. -0.01 -0.01 0.00
Coreference Resolution 0.00 0.00 0.00
Dialogue Generation -0.02 -0.01 0.00
Fill in The Blank 0.04 0.04 0.04
Information Extr. 0.03 0.06 0.04
Named Entity Rec. 0.01 0.02 0.01
Program Execution -0.02 -0.01 -0.03
Question Answering 0.14 0.13 0.13
Question Generation 0.06 0.06 0.06
Question Rewriting 0.03 0.04 0.04
Question Underst. -0.01 -0.03 -0.02
Sentence Composition 0.10 0.11 0.11
Sentiment Analysis 0.03 0.03 0.03
Summarization 0.08 0.08 0.08
Text Categorization -0.03 -0.02 -0.02
Text Completion 0.01 0.02 0.00
Text Matching 0.00 0.01 0.01
Text Quality Eval. 0.00 -0.01 -0.01
Text to Code -0.02 -0.01 0.01
Textual Entailment 0.03 0.01 0.03
Title Generation 0.01 0.01 0.01
Word Semantics -0.02 -0.03 -0.01
Wrong Candidate Gen. 0.01 0.03 0.00
Overall 0.02 0.02 0.02
Table 2: The average Pearson correlations between the number of tokens (Tok.) and task perf., word position deviation (Pos.) and task perf., and lexical deviation (Lex.) and task perf. for LLaMA 3 70B. Each task family has five downstream tasks. The average p-value is 0.05.
Task Family Min Max Avg ($\downarrow$)
Title Generation 24 53 36.0
Fill in The Blank 25 60 43.2
Summarization 21 123 52.0
Text Quality Eval. 42 67 52.6
Named Entity Rec. 38 77 52.8
Word Semantics 39 74 53.8
Sentiment Analysis 24 115 56.6
Sentence Composition 19 84 61.8
Answerability Class. 47 70 63.6
Text Completion 40 111 66.0
Dialogue Generation 27 186 67.8
Question Answering 45 106 72.0
Program Execution 43 110 75.4
Coreference Resolution 24 219 78.8
Information Extr. 70 91 79.0
Textual Entailment 61 111 81.2
Text Categorization 55 116 88.2
Text Matching 34 203 91.0
Question Generation 29 160 102.2
Wrong Candidate Gen. 105 192 129.4
Commonsense Class. 49 202 148.8
Question Rewriting 46 753 216.0
Question Understanding 125 581 291.8
Text to Code 121 638 390.6
Overall 19 753 102.1
Table 3: The number of tokens per input prompt (computed using the GPT-2 tokenizer). Min shows the shortest prompt, Max the longest prompt, and Avg the average prompt length for all five tasks.
Appendix C Additional Results
Prompt Complexity Correlations. Table 2 shows the individual correlations between the three markers of prompt complexity and downstream task performance as described in Section B.2. All correlations are close to zero, and only very small correlations exist for question answering (0.13 - 0.14) and sentence composition (0.10 - 0.11) with $p<0.05$. We could generally find no evidence that the complexity of the prompt contributes to gains or losses in performance outcomes.
Prompt Lengths. We conducted a post-hoc analysis of the number of input tokens across various task families using the GPT-2 tokenizer. As shown in Table 3, the average number of tokens varies significantly depending on the nature of the task, ranging from 36 tokens for title generation tasks to 390 tokens for text-to-code tasks. The overall distribution of input tokens is even broader when looking at individual tasks, with the shortest task, sentence composition, using as few as 19 tokens, and the longest, question rewriting, reaching up to 753 tokens. This wide spectrum of prompt lengths indicates that our study covers a diverse array of real-world use cases, from concise to extended prompts. By analyzing 120 tasks across 24 task families, we believe our conclusions hold across a broad range of prompt lengths, including those much shorter (e.g., fewer than 36 tokens) and much longer (e.g., more than 800 tokens).
Results by Task and Type. We show a detailed table of avgerage downstream task performance gain or loss over the original prompt of a task when paraphrasing the prompt using modifications of one of the six different paraphrase groups in Table 10. We decompose this table for each model in Tables 12, 11, 13, 15 and 14.
Maximum Potential Gain. To understand how much models would possibly benefit from adjusting prompts in optimal ways as explained in Table 8, we also show the maximum potential gain in Figure 9 of the teaser on the first page of this paper (Figure 1). This gives an upper bound to what is possible if prompts were optimally adjusted with paraphrase types. Again, we note that although finding these optimal prompt adjustments is challenging, this task is not impossible and gives space for future work.
Model Scale Comparison. Figure 12 compares model size between LLaMA 3 8B and LLaMA 3 70B. Observe how the smaller 8B parameter model benefits more across tasks when compared to its larger 70B version. Only for program execution, information extraction, sentence completion, question rewriting, and question generation, the 70B model shows a larger gain.
Upon closer examination, the performance difference between the two models across various tasks highlights several trends. The smaller 8B model consistently outperforms the 70B model in tasks such as text categorization, text quality evaluation, named entity recognition, and sentiment analysis. This suggests that for these specific tasks, the architectural or training enhancements in the 8B model may be more efficiently leveraged, potentially due to better optimization or more effective use of parameter space.
In contrast, the 70B model demonstrates improvements over the 8B model in tasks like program execution, information extraction, and sentence completion. This indicates that for these more complex or nuanced tasks, the increased parameter count of the 70B model likely provides a richer representation space, allowing it to capture and utilize more intricate patterns and dependencies in the data.
Figure 11: The average percentage in downstream task performance gain or loss for applying different paraphrase types across all five models. This decomposes the groups of Figure 3 into their individual types.
Figure 12: A comparison between model scale for LLaMA 3 8B (blue) and 70B (orange) for different tasks.
Task ID Dataset Name Domain Inp. Tokens Out. Tokens URL
Text Completion - Morphology (+11.4% $\uparrow$), Syntax (-5.3% $\downarrow$)
   105 ROCStories Story, Commonsense 39.3 9.0 Link
   297 ROCStories Story 56.1 1.0 Link
   268 CaseHold Law 269.1 1.0 Link
   138 Detoxifying LMs Social Media 59.6 2.0 Link
   296 ROCStories Story 56.1 1.0 Link
Question Answering - Extremes (+24.8% $\uparrow$), Lexicon (-3.0% $\downarrow$)
   1441 DoQA Movies Movies, Dialogue 190.6 16.0 Link
   2 Quoref Wikipedia 356.6 1.7 Link
   24 Cosmos QA Personal Narratives 82.5 8.3 Link
   332 TellMeWhy Story 50.4 8.9 Link
   310 RACE English Exams 351.6 1.0 Link
Fill in The Blank - Extremes (+18.4% $\uparrow$), Discourse (-5.0% $\downarrow$)
   672 NumerSense Commonsense 11.1 1.0 Link
   277 StereoSet Stereotypes 8.8 1.0 Link
   965 LibriSpeech ASR Books 42.1 1.0 Link
   1217 ATOMIC Sociology, Commonsense 5.3 1.1 Link
   1360 NumerSense Concepts and Relations 26.1 1.0 Link
Sentiment Analysis - Morphology (+15.6% $\uparrow$), Discourse (-2.1% $\downarrow$)
   420 PerSenT News 352.6 1.0 Link
   889 GoEmotions Narrative, Dialogue 12.3 1.0 Link
   1312 Amazon Review Polarity Reviews 56.7 1.0 Link
   1535 Daily Dialog Dialogue 137.8 1.0 Link
   284 IMDB Movie Reviews 229.9 1.0 Link
Program Execution - Morphology (+1.9% $\uparrow$), Discourse (-4.2% $\downarrow$)
   208 Combinations of List Mathematics 5.0 22.1 -
   378 Reverse Words Image Captions 21.4 10.4 -
   1148 Maximum ASCII Value Computer Science 1.0 1.0 -
   94 Calculate Mean Code 6.0 1.0 Link
   99 Reverse Elements Code 17.9 6.3 -
Question Generation - Morphology (+21.5% $\uparrow$), Discourse (+1.3% $\uparrow$)
   6 MCTACO News, Wiki, Law, History 18.9 8.2 Link
   1602 WebQuestions Knowledge Base 9.6 6.6 Link
   599 CUAD Law 3153.9 43.3 Link
   405 NarrativeQA Books, Movies 575.7 8.6 Link
   739 LhoestQ Web 142.1 10.4 -
Table 4: Overview of the tasks in this study with their average input and output tokens across all examples. We also show the two paraphrase types with the highest and lowest gain or loss across all five sampled tasks within a task category. Table 1/4.
Task ID Dataset Name Domain Inp. Tokens Out. Tokens URL
Text to Code - Morphology (+2.4% $\uparrow$), Lexico-Syntax (-5.5% $\downarrow$)
   869 CFQ MCD1 SQL 41.2 1.0 Link
   956 LeetCode 420 Mathematics 3.0 1.0 Link
   128 SCAN Machine Learning 7.0 10.8 Link
   107 SPLASH SQL 13.3 17.5 Link
   211 Logic2Text Wikipedia, Logic 52.6 1.0 Link
Question Rewriting - Morphology (-4.7% $\downarrow$), Syntax (-17.2% $\downarrow$)
   402 GrailQA Knowledge Base 125.5 11.2 Link
   671 AmbigQA Wikipedia 8.9 14.4 Link
   34 WinoGrande Commonsense 26.1 20.0 Link
   1622 Disfl QA Wikipedia 14.6 9.9 Link
   670 AmbigQA Wikipedia 8.9 12.1 Link
Summarization - Lexicon (+10.8% $\uparrow$), Syntax (+2.6% $\uparrow$)
   1290 XSum News 371.8 21.1 Link
   1357 XLSum News 459.2 22.1 Link
   668 SciTLDR Scientific Papers 159.5 20.6 Link
   589 Amazon Food Reviews Reviews 79.6 4.2 Link
   672 Amazon / Yelp Summ. Reviews 404.0 50.7 Link
Commonsense Classification - Morphology (+14.6% $\uparrow$), Lexicon (-2.7% $\downarrow$)
   116 Com2Sense Concepts and Relations 19.2 1.0 Link
   1204 ATOMIC Social Commonsense 10.4 1.0 Link
   1209 ATOMIC Physical Commonsense 7.0 1.0 Link
   1199 ATOMIC Social Commonsense 7.4 1.0 Link
   1208 ATOMIC Social Commonsense 7.0 1.0 Link
Text Matching - Syntax (+5.4% $\uparrow$), Discourse (-12.2% $\downarrow$)
   1288 GLUE MRPC News, Web 45.9 1.0 Link
   276 Enhanced WSC Dialogue, Narrative 39.8 1.0 Link
   910 Bianet News 48.5 1.0 Link
   148 AFS Government and Politics 25.6 1.0 Link
   624 OHSUMED Scientific Papers 188.5 11.2 Link
Word Semantics - Morphology (+14.9% $\uparrow$), Lexico-Syntax (-6.5% $\downarrow$)
   1585 ROOT09 Misc. 1.0 1.0 Link
   1582 BLESS Misc. 1.0 1.0 Link
   141 Odd-Man-Out Card Game 9.0 1.1 Link
   142 Odd-Man-Out Card Game 5.3 1.1 Link
   458 MATRES News 49.2 1.0 Link
Table 5: Overview of the tasks in this study with their average input and output tokens across all examples. We also show the two paraphrase types with the highest and lowest gain or loss across all five sampled tasks within a task category. Table 2/4.
Task ID Dataset Name Domain Inp. Tokens Out. Tokens URL
Question Understanding - Morphology (+15.1% $\uparrow$), Discourse (-4.1% $\uparrow$)
   19 MCTACO News, Wiki, Law, History 27.8 1.0 Link
   46 Misc. Pop, Nat. Science, History 14.8 1.0 -
   18 MCTACO News, Wiki, Law, History 24.9 1.0 Link
   1289 TREC Misc. 10.2 1.0 Link
Text Quality Evaluation - Morphology (+7.7% $\uparrow$), Discourse (-8.6% $\uparrow$)
   616 CoLA Linguistics 7.9 1.0 Link
   1284 Human Ratings of NLG Dialogue, Restaurants 23.8 1.0 Link
   1623 Disfl QA Wikipedia 12.2 1.0 Link
   1283 Human Ratings of NLG Dialogue, Restaurants 24.9 1.0 Link
   1341 MSR Text Compression News, Dialogue, Misc. 19.4 1.0 Link
Dialogue Generation - Morphology (+9.3% $\uparrow$), Discourse (-3.9% $\uparrow$)
   639 MultiWOZ v2.2 Dialogue 13.2 12.6 Link
   1730 PersonaChat Dialogue 143.0 9.9 Link
   576 Curiosity Dialogs Concepts and Relations 66.2 23.0 Link
   361 Spolin Dialogue 32.3 1.0 Link
   1603 SMCalFlow Dialogue 8.0 8.7 Link
Coreference Resolution - Others (+25.1% $\uparrow$), Lexico-Syntax (+1.6% $\uparrow$)
   1391 WinoGrande Physical Commonsense 22.8 1.0 Link
   648 Winograd WSC Narrative 19.4 1.7 Link
   133 WinoWhy Concepts and Relations 43.8 1.0 Link
   330 GAP Wikipedia 73.8 1.4 Link
   329 GAP Wikipedia 81.2 1.0 Link
Answerability Classification - Lexico-Syntax (+11.8% $\uparrow$), Others (-3.5% $\uparrow$)
   290 TellMeWhy Story 51.0 1.3 Link
   50 MultiRC News, Wiki, Law, History 23.3 1.0 Link
   349 SQuAD 2.0 Wikipedia 140.5 1.0 Link
   20 MCTACO News, Wiki, Law, History 25.0 1.0 Link
   1640 Adversarial QA Wikipedia 127.3 1.0 Link
Wrong Candidate Generation - Morphology (+26.0% $\uparrow$), Syntax (-1.4% $\uparrow$)
   135 Winowhy Concepts and Relations 26.0 12.5 Link
   42 QASC Nat. Science 22.5 1.6 Link
   55 MultiRC News, Wiki, Law, History 308.0 3.4 Link
   11 MCTACO News, Wiki, Law, History 27.6 4.9 Link
   631 DBPedia 14 Wikipedia 53.4 1.4 Link
Table 6: Overview of the tasks in this study with their average input and output tokens across all examples. We also show the two paraphrase types with the highest and lowest gain or loss across all five sampled tasks within a task category. Table 3/4.
Task ID Dataset Name Domain Inp. Tokens Out. Tokens URL
Sentence Composition - Discourse (+3.8% $\uparrow$), Morphology (-5.0% $\downarrow$)
   184 SNLI Image Captions 23.4 8.3 Link
   1613 SICK Image & Video Captions 12.6 9.2 Link
   1530 SciTailv1.1 Nat. Science 19.6 12.2 Link
   1368 HealthFact Healthcare 566.0 11.2 Link
   1364 HANS Movie Reviews 8.6 6.0 Link
Textual Entailment - Morphology (+17.5% $\uparrow$), Discourse (-6.2% $\downarrow$)
   640 e-SNLI Misc. 23.1 1.0 Link
   201 MultiNLI History, Fiction, Dialogue, Law 54.9 1.0 Link
   1387 ANLI R3 Misc. 67.8 1.0 Link
   190 SNLI Image Captions 24.2 1.0 Link
   1612 SICK Image & Video Captions 21.1 1.0 Link
Named Entity Recognition - Morphology (+9.2% $\uparrow$), Others (-15.3% $\downarrow$)
   1480 JNLPBA Bioinformatics 29.5 2.5 Link
   1486 ANEM Clinical Knowledge 30.3 1.2 Link
   959 E2E Restaurants 27.1 2.1 Link
   1483 ChemProt Chemistry 14.8 1.2 Link
   1481 BC2GM Bioinformatics 30.5 2.4 Link
Text Categorization - Morphology (+9.2% $\uparrow$), Discourse (-7.8% $\downarrow$)
   1495 ADE Corpus V2 Clinical Knowledge, Healthcare 17.7 3.0 Link
   1489 Sarcasm in Twitter Social Media 17.7 1.0 Link
   1308 Amazon Reviews Reviews 64.2 1.0 Link
   617 Amazon Reviews Reviews 68.0 1.3 Link
   1541 AG News News 37.8 1.0 Link
Title Generation - Morphology (+12.1% $\uparrow$), Lexico-Syntax (+0.9% $\uparrow$)
   220 ROCStories Narrative, Story 60.0 1.0 Link
   288 GigaWord News 29.9 8.8 Link
   619 OHSUMED Reviews 157.5 11.2 Link
   1540 PeerRead Computer Science 164.9 8.4 Link
   1659 BillSum Government and Politics 180.3 18.4 Link
Information Extraction - Morphology (+7.5% $\uparrow$), Others (-5.3% $\downarrow$)
   646 Winograd WSC Narrative 16.4 1.7 Link
   1413 DART Wikipedia 8.9 1.7 Link
   748 GLUCOSE Story 55.4 11.5 Link
   1411 DART Wikipedia 9.1 2.2 Link
   1506 Synthetic Pop 22.5 3.0 -
Table 7: Overview of the tasks in this study with their average input and output tokens across all examples. We also show the two paraphrase types with the highest and lowest gain or loss across all five sampled tasks within a task category. Table 4/4.
Paraphrase Group Paraphrase Type Model Output ROUGE-L
Morphology Inflectional Great Value 1.00
Extremes Entailment Great Value 1.00
Syntax Negation Switching Great Value Cat Food 0.67
Morphology Modal Verb Great Value Cat Food 0.67
Morphology Derivational Great Value Cat Food 0.67
Discourse Direct/Indirect Style Altern. Great Value Cat Food 0.67
Lexico-Syntax Synthetic/Analytic Sub. Great Value for Cats 0.67
Lexico-Syntax Opp. Pol. Sub. (Habitual) Great Value for Multiple Cats 0.57
Lexicon Spelling Great Value for Multiple Cats 0.57
Extremes Non-Paraphrase Good Value Cat Food 0.33
Original Original Good Value Cat Food 0.33
Discourse Syntax/Discourse Structure Good Value for Cats 0.33
Syntax Subordination and Nesting Purrfect Kitten Food 0.00
Others Addition/Deletion Kittens loved the food. 0.00
Discourse Punctuation Kittens loved the food 0.00
Others Change of Order Cats love the food. 0.00
Extremes Identity Cats loved the food 0.00
Syntax Coordination Kittens Loved the Food 0.00
Syntax Ellipsis Good for Kittens 0.00
Syntax Diathesis Alternation Positive 0.00
Lexico-Syntax Opp. Pol. Sub. (Contextual) Good Quality Cat Food 0.00
Lexico-Syntax Converse Sub. Cats Loved Food 0.00
Lexicon Same Pol. Sub. (Named Ent.) Kittens loved the food. 0.00
Lexicon Same Pol. Sub. (Habitual) Perfect for Multiple Cats 0.00
Lexicon Same Pol. Sub. (Contextual) Multiple Cat Food 0.00
Lexicon Change of Format Cats loved the food 0.00
Others Semantic-Based Cats loved the food 0.00
Table 8: How we calculate the potential performance gain for paraphrasing prompts using a selected example from the Amazon food review summarization task McAuley and Leskovec (2013) with LLaMA 3 8B. Median Gain. To calculate the median gain (as in Figure 1), we consider all paraphrases of the prompt that are better than the original prompt (green and yellow). Median Loss. To calculate the median loss (as in Figure 1), we consider all paraphrases of the prompt that are worse than the original prompt (red). Max. We calculate the maximum possible gain by aggregating only the best performing paraphrases of the prompts highlighted in yellow.
Paraphrase Type # Examples Computed
Morphology 360k
   Derivational Changes
   Inflectional Changes
   Modal Verb Changes
Lexicon 600k
   Spelling changes
   Change of format
   Same Polarity Substitution (contextual)
   Same Polarity Substitution (habitual)
   Same Polarity Substitution (named ent.)
Lexico-syntactic 480k
   Converse substitution
   Opposite polarity substitution (contextual)
   Opposite polarity substitution (habitual)
   Synthetic/analytic substitution
Syntax 600k
   Coordination changes
   Diathesis alternation
   Ellipsis
   Negation switching
   Subordination and nesting changes
Discourse 360k
   Direct/indirect style alternations
   Punctuation changes
   Syntax/discourse structure changes
Others 720k
   Addition/Deletion
   Change of order
   Semantic-based
   Entailment
   Identity
   Non-paraphrase
Total 3,24m
Table 9: An overview of the considered types of paraphrase prompts, categorized into their six main groups. The numbers indicate how many task examples with instructions of that paraphrase type group were run in our experiments (this is a product of all models, tasks, and number of examples per task). The number of paraphrases within a group is equally balanced (i.e., derivational changes have occurred the same amount of time as inflectional changes).
Task Family Morphology Syntax Lexicon Lex.-Syn. Discourse Others
Answerability Classification +7.0% -3.0% +6.2% +11.8% +0.9% -3.5%
Commonsense Classification +14.6% +2.9% -2.7% -1.9% +0.6% -1.6%
Coreference Resolution +8.3% +9.4% +4.6% +1.6% +3.3% +25.1%
Dialogue Generation +9.3% +3.4% +6.3% +9.2% -3.9% +4.6%
Fill in The Blank +10.0% -2.0% +4.7% +2.7% -5.0% +18.4%
Information Extraction +7.5% -1.7% -1.0% -5.2% +6.1% -5.3%
Named Entity Recognition +9.2% -8.7% -2.1% +2.0% +8.7% -15.3%
Program Execution +1.94% -3.1% -1.5% -1.7% -4.2% -2.8%
Question Answering +10.3% +1.8% -3.0% +9.9% +5.0% +24.8%
Question Generation +21.5% +5.9% +2.5% +5.9% +1.3% +4.1%
Question Rewriting -4.7% -17.2% -6.6% -5.8% -8.2% -8.6%
Question Understanding +15.1% +3.5% -2.7% -2.8% -4.1% -3.3%
Sentence Composition -5.0% +3.4% -2.9% -2.7% +3.8% +2.9%
Sentiment Analysis +15.6% +1.6% +1.7% +1.9% -2.1% -1.7%
Summarization +6.9% +2.6% +10.8% +4.8% +7.8% +7.5%
Text Categorization +9.2% -5.3% -7.0% -1.6% -7.8% -2.4%
Text Completion +11.4% -5.3% -3.0% -2.6% -3.1% +1.7%
Text Matching +0.5% +5.4% -1.1% +0.3% -12.2% -1.9%
Text Quality Evaluation +7.7% -1.1% +1.4% +1.2% -8.6% +1.4%
Text to Code +2.4% -2.1% -1.7% -5.5% -0.6% +0.7%
Textual Entailment +17.5% -4.2% +5.2% +5.3% -6.2% -4.6%
Title Generation +12.1% +5.8% +7.1% +1.7% +0.9% +7.4%
Word Semantics +14.9% -0.5% -2.7% -6.5% +3.1% +2.6%
Wrong Candidate Generation +26.0% -1.4% +8.9% +7.5% -2.9% +5.9%
Table 10: The average downstream task performance gain or loss over the original prompt when paraphrased with a specific type from one of the six groups (columns) of tasks within a certain task family (rows) as an average over all models. We calculate the average across all paraphrase types in one of the six paraphrase groups and across all five tasks within one of the 24 categories. Bold indicates the highest score per column. Small changes between -3% and +3% are not colored.
Task Family Morphology Syntax Lexicon Lex.-Syn. Discourse Others
Answerability Classification -9.0% $\downarrow$ -2.1% +21.9% $\uparrow$ +57.6% $\uparrow$ 0.0% 0.0%
Commonsense Classification +6.4% $\uparrow$ 0.0% -3.1% $\downarrow$ -2.4% 0.0% 0.0%
Coreference Resolution +6.4% $\uparrow$ +35.0% $\uparrow$ +10.8% $\uparrow$ +18.2% $\uparrow$ 0.0% +25.0% $\uparrow$
Dialogue Generation +8.4% $\uparrow$ +1.6% +4.5% $\uparrow$ +1.6% -2.0% +28.6% $\uparrow$
Fill in The Blank -18.4% $\downarrow$ +4.5% $\uparrow$ +15.4% $\uparrow$ +3.3% $\uparrow$ -10.4% $\downarrow$ +36.0% $\uparrow$
Information Extraction +8.7% $\uparrow$ -11.4% $\downarrow$ -13.6% $\downarrow$ -15.3% $\downarrow$ +17.9% $\uparrow$ -20.0% $\downarrow$
Named Entity Recognition +9.0% $\uparrow$ -42.3% $\downarrow$ -16.2% $\downarrow$ +20.6% $\uparrow$ +49.7% $\uparrow$ -71.6% $\downarrow$
Program Execution +9.7% $\uparrow$ -11.7% $\downarrow$ -7.9% $\downarrow$ -7.5% $\downarrow$ -18.9% $\downarrow$ -23.7% $\downarrow$
Question Answering +7.6% $\uparrow$ +9.6% $\uparrow$ -6.6% $\downarrow$ +4.4% $\uparrow$ -4.9% $\downarrow$ +13.0% $\uparrow$
Question Generation +19.4% $\uparrow$ -16.4% $\downarrow$ -10.8% $\downarrow$ -11.8% $\downarrow$ -10.7% $\downarrow$ -6.5% $\downarrow$
Question Rewriting -8.3% $\downarrow$ -9.0% $\downarrow$ -5.8% $\downarrow$ -8.1% $\downarrow$ -26.1% $\downarrow$ -15.9% $\downarrow$
Question Understanding -12.3% $\downarrow$ -14.6% $\downarrow$ -17.5% $\downarrow$ -21.3% $\downarrow$ -13.2% $\downarrow$ -34.8% $\downarrow$
Sentence Composition -23.7% $\downarrow$ +6.3% $\uparrow$ +4.6% $\uparrow$ +4.7% $\uparrow$ +6.4% $\uparrow$ +1.7%
Sentiment Analysis +3.5% $\uparrow$ 0.0% -2.5% -5.6% $\downarrow$ 0.0% 0.0%
Summarization +14.6% $\uparrow$ -4.0% $\downarrow$ +7.9% $\uparrow$ +9.5% $\uparrow$ -5.3% $\downarrow$ +4.3% $\uparrow$
Text Categorization +5.3% $\uparrow$ -19.5% $\downarrow$ -30.8% $\downarrow$ -13.2% $\downarrow$ -41.2% $\downarrow$ 0.0%
Text Completion +20.0% $\uparrow$ -33.7% $\downarrow$ -30.3% $\downarrow$ -18.8% $\downarrow$ -17.1% $\downarrow$ -10.0% $\downarrow$
Text Matching +3.0% +22.1% $\uparrow$ +5.5% $\uparrow$ -1.0% -51.3% $\downarrow$ -22.7% $\downarrow$
Text Quality Evaluation -5.0% $\downarrow$ -3.2% $\downarrow$ -4.0% $\downarrow$ -7.4% $\downarrow$ -65.4% $\downarrow$ 0.0%
Text to Code +8.0% $\uparrow$ +3.9% $\uparrow$ +10.0% $\uparrow$ -2.6% -1.0% 0.0%
Textual Entailment +16.3% $\uparrow$ -31.1% $\downarrow$ +17.0% $\uparrow$ +14.9% $\uparrow$ -35.5% $\downarrow$ -1.2%
Title Generation +11.6% $\uparrow$ +9.9% $\uparrow$ +9.3% $\uparrow$ -13.1% $\downarrow$ -6.3% $\downarrow$ -1.8%
Word Semantics +4.9% $\uparrow$ -7.8% $\downarrow$ -7.7% $\downarrow$ -19.6% $\downarrow$ +9.8% $\uparrow$ -0.1%
Wrong Candidate Generation +24.5% $\uparrow$ +6.4% $\uparrow$ +8.4% $\uparrow$ +9.6% $\uparrow$ +8.9% $\uparrow$ +5.1% $\uparrow$
Table 11: The average downstream task performance gain or loss over the original prompt when paraphrased with a specific type from one of the six groups (columns) of tasks within a certain task family (rows) for Mixtral 8x7B Instruct (47B). We calculate the average across all paraphrase types in one of the six paraphrase groups and across all five tasks within one of the 24 categories. Bold indicates the highest score per column. Small changes between -3% and +3% are not colored.
Task Family Morphology Syntax Lexicon Lex.-Syn. Discourse Others
Answerability Classification +34.7% $\uparrow$ -10.5% $\downarrow$ +9.2% $\uparrow$ +5.0% $\uparrow$ +4.6% $\uparrow$ -17.6% $\downarrow$
Commonsense Classification +44.9% $\uparrow$ +14.3% $\uparrow$ -10.4% $\downarrow$ +11.2% $\uparrow$ +2.8% -7.9% $\downarrow$
Coreference Resolution +23.1% $\uparrow$ -10.0% $\downarrow$ +11.1% $\uparrow$ -8.9% $\downarrow$ -4.8% $\downarrow$ +0.4%
Dialogue Generation +15.5% $\uparrow$ -11.3% $\downarrow$ +11.4% $\uparrow$ +11.8% $\uparrow$ +0.9% -8.9% $\downarrow$
Fill in The Blank +30.7% $\uparrow$ -11.3% $\downarrow$ +12.5% $\uparrow$ -7.9% $\downarrow$ -16.7% $\downarrow$ +17.6% $\uparrow$
Information Extraction +17.8% $\uparrow$ -4.7% $\downarrow$ +8.2% $\uparrow$ -8.3% $\downarrow$ +12.6% $\uparrow$ -6.5% $\downarrow$
Named Entity Recognition +17.2% $\uparrow$ -5.2% $\downarrow$ +9.0% $\uparrow$ -9.5% $\downarrow$ -5.7% $\downarrow$ -5.1% $\downarrow$
Program Execution +22.6% $\uparrow$ -7.1% $\downarrow$ +8.3% $\uparrow$ -9.1% $\downarrow$ +4.7% $\uparrow$ +9.6% $\uparrow$
Question Answering +17.6% $\uparrow$ +4.7% $\uparrow$ +15.9% $\uparrow$ +8.0% $\uparrow$ +12.4% $\uparrow$ -8.3% $\downarrow$
Question Generation +18.1% $\uparrow$ +11.1% $\uparrow$ +14.0% $\uparrow$ +16.6% $\uparrow$ -8.7% $\downarrow$ +30.4% $\uparrow$
Question Rewriting -13.6% $\downarrow$ -4.0% $\downarrow$ -12.3% $\downarrow$ +6.8% $\uparrow$ -5.9% $\downarrow$ -7.5% $\downarrow$
Question Understanding +35.3% $\uparrow$ +17.7% $\uparrow$ -11.3% $\downarrow$ -10.6% $\downarrow$ -9.6% $\downarrow$ -2.3%
Sentence Composition +14.7% $\uparrow$ -6.2% $\downarrow$ -9.4% $\downarrow$ +6.5% $\uparrow$ +6.8% $\uparrow$ -4.5% $\downarrow$
Sentiment Analysis +30.6% $\uparrow$ +8.1% $\uparrow$ +10.0% $\uparrow$ +10.1% $\uparrow$ -9.5% $\downarrow$ -8.5% $\downarrow$
Summarization -16.1% $\downarrow$ -9.9% $\downarrow$ +11.3% $\uparrow$ +10.1% $\uparrow$ +16.9% $\uparrow$ +10.9% $\uparrow$
Text Categorization +26.3% $\uparrow$ -10.9% $\downarrow$ -9.5% $\downarrow$ +8.0% $\uparrow$ -8.0% $\downarrow$ -12.1% $\downarrow$
Text Completion +24.0% $\uparrow$ +11.7% $\uparrow$ +14.3% $\uparrow$ +10.4% $\uparrow$ -4.5% $\downarrow$ +8.8% $\uparrow$
Text Matching -10.5% $\downarrow$ -4.4% $\downarrow$ -6.8% $\downarrow$ +4.1% $\uparrow$ -9.9% $\downarrow$ +13.8% $\uparrow$
Text Quality Evaluation +28.2% $\uparrow$ -7.3% $\downarrow$ +11.6% $\uparrow$ +6.7% $\uparrow$ +22.5% $\uparrow$ +7.1% $\uparrow$
Text to Code -11.6% $\downarrow$ -14.3% $\downarrow$ -11.4% $\downarrow$ -19.2% $\downarrow$ +10.8% $\uparrow$ +1.2%
Textual Entailment +43.2% $\uparrow$ +9.1% $\uparrow$ +13.8% $\uparrow$ +11.8% $\uparrow$ +4.9% $\uparrow$ -23.3% $\downarrow$
Title Generation +22.3% $\uparrow$ +9.6% $\uparrow$ +12.0% $\uparrow$ +7.7% $\uparrow$ -9.4% $\downarrow$ -7.3% $\downarrow$
Word Semantics +38.5% $\uparrow$ +5.9% $\uparrow$ -9.0% $\downarrow$ -12.9% $\downarrow$ +1.7% +8.3% $\uparrow$
Wrong Candidate Generation +18.9% $\uparrow$ -7.8% $\downarrow$ +9.6% $\uparrow$ +8.2% $\uparrow$ -6.1% $\downarrow$ +6.7% $\uparrow$
Table 12: The average downstream task performance gain or loss over the original prompt when paraphrased with a specific type from one of the six groups (columns) of tasks within a certain task family (rows) for Gemma 7B Instruct (7B). We calculate the average across all paraphrase types in one of the six paraphrase groups and across all five tasks within one of the 24 categories. Bold indicatess the highest score per paraphrase group (column). Small changes between -3% and +3% are not colored.
Task Family Morphology Syntax Lexicon Lex.-Syn. Discourse Others
Answerability Classification +3.2% $\uparrow$ -2.6% -0.1% -1.7% 0.0% 0.0%
Commonsense Classification +10.4% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Coreference Resolution +5.8% $\uparrow$ 0.0% -1.3% 0.0% 0.0% 0.0%
Dialogue Generation +11.2% $\uparrow$ +19.4% $\uparrow$ +5.6% $\uparrow$ +18.4% $\uparrow$ -8.5% $\downarrow$ 0.0%
Fill in The Blank +13.3% $\uparrow$ -3.6% $\downarrow$ 0.0% 0.0% 0.0% 0.0%
Information Extraction +5.5% $\uparrow$ +6.6% $\uparrow$ +3.9% $\uparrow$ -2.8% -0.4% 0.0%
Named Entity Recognition +11.5% $\uparrow$ 0.0% -3.3% $\downarrow$ -0.7% 0.0% 0.0%
Program Execution -9.9% $\downarrow$ -4.7% $\downarrow$ -3.5% $\downarrow$ -3.2% $\downarrow$ -5.9% $\downarrow$ 0.0%
Question Answering +10.6% $\uparrow$ -10.7% $\downarrow$ -6.8% $\downarrow$ +16.2% $\uparrow$ +8.3% $\uparrow$ +30.0% $\uparrow$
Question Generation +38.3% $\uparrow$ +13.4% $\uparrow$ +13.1% $\uparrow$ -8.9% $\downarrow$ +7.3% $\uparrow$ +12.7% $\uparrow$
Question Rewriting -6.5% $\downarrow$ -9.1% $\downarrow$ -6.9% $\downarrow$ -5.9% $\downarrow$ -2.4% -4.2% $\downarrow$
Question Understanding +10.9% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Sentence Composition -14.5% $\downarrow$ +2.1% +4.8% $\uparrow$ -4.8% $\downarrow$ -0.7% +3.8% $\uparrow$
Sentiment Analysis +7.5% $\uparrow$ 0.0% -1.2% 0.0% 0.0% 0.0%
Summarization +23.6% $\uparrow$ +6.6% $\uparrow$ +15.5% $\uparrow$ -9.2% $\downarrow$ +15.7% $\uparrow$ -11.1% $\downarrow$
Text Categorization +3.6% $\uparrow$ 0.0% -0.2% 0.0% 0.0% 0.0%
Text Completion +5.5% $\uparrow$ +1.6% +2.7% +3.3% $\uparrow$ -0.4% -0.1%
Text Matching +3.5% $\uparrow$ 0.0% 0.0% -0.2% 0.0% 0.0%
Text Quality Evaluation +6.0% $\uparrow$ 0.0% -0.4% 0.0% 0.0% 0.0%
Text to Code +4.8% $\uparrow$ -0.7% -1.4% -1.0% -4.3% $\downarrow$ 0.0%
Textual Entailment +11.0% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Title Generation +12.8% $\uparrow$ +8.2% $\uparrow$ +14.2% $\uparrow$ +12.0% $\uparrow$ +5.6% $\uparrow$ +10.5% $\uparrow$
Word Semantics +14.1% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Wrong Candidate Generation +37.8% $\uparrow$ -3.1% $\downarrow$ +10.2% $\uparrow$ +10.9% $\uparrow$ -11.7% $\downarrow$ -2.0%
Table 13: The average downstream task performance gain or loss over the original prompt when paraphrased with a specific type from one of the six groups (columns) of tasks within a certain task family (rows) for Command R+ (104B). We calculate the average across all paraphrase types in one of the six paraphrase groups and across all five tasks within one of the 24 categories. Bold indicatess the highest score per paraphrase group (column). Small changes between -3% and +3% are not colored.
Task Family Morphology Syntax Lexicon Lex.-Syn. Discourse Others
Answerability Classification +4.2% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Commonsense Classification +7.8% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Coreference Resolution +8.0% $\uparrow$ 0.0% +2.2% -0.7% 0.0% 0.0%
Dialogue Generation +6.9% $\uparrow$ +4.9% $\uparrow$ +6.0% $\uparrow$ +7.7% $\uparrow$ -9.6% $\downarrow$ +3.3% $\uparrow$
Fill in The Blank +19.7% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Information Extraction +3.1% $\uparrow$ -0.6% -2.4% +1.2% -0.5% 0.0%
Named Entity Recognition +5.8% $\uparrow$ -1.4% +0.6% -0.2% 0.0% 0.0%
Program Execution -8.5% $\downarrow$ +5.5% $\uparrow$ +7.3% $\uparrow$ +6.1% $\uparrow$ -1.1% 0.0%
Question Answering +11.5% $\uparrow$ +5.4% $\uparrow$ +9.5% $\uparrow$ +10.7% $\uparrow$ -0.6% +37.7% $\uparrow$
Question Generation +19.4% $\uparrow$ +13.6% $\uparrow$ +14.1% $\uparrow$ +15.7% $\uparrow$ +6.5% $\uparrow$ -15.0% $\downarrow$
Question Rewriting +9.1% $\uparrow$ -10.0% $\downarrow$ -9.2% $\downarrow$ -6.6% $\downarrow$ -7.5% $\downarrow$ -8.2% $\downarrow$
Question Understanding +9.6% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Sentence Composition +7.3% $\uparrow$ +10.3% $\uparrow$ -7.9% $\downarrow$ -8.5% $\downarrow$ +11.0% $\uparrow$ -6.8% $\downarrow$
Sentiment Analysis +4.2% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Summarization -16.4% $\downarrow$ +10.4% $\uparrow$ +11.2% $\uparrow$ +12.7% $\uparrow$ +9.5% $\uparrow$ +30.9% $\uparrow$
Text Categorization +5.0% $\uparrow$ 0.0% -0.1% 0.0% 0.0% 0.0%
Text Completion +4.9% $\uparrow$ +15.9% $\uparrow$ -4.3% $\downarrow$ -7.1% $\downarrow$ 0.0% 0.0%
Text Matching +4.1% $\uparrow$ 0.0% -0.3% 0.0% 0.0% 0.0%
Text Quality Evaluation +5.3% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Text to Code +10.1% $\uparrow$ -4.1% $\downarrow$ -3.8% $\downarrow$ -3.2% $\downarrow$ -6.5% $\downarrow$ +4.5% $\uparrow$
Textual Entailment +13.4% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Title Generation +10.9% $\uparrow$ +13.1% $\uparrow$ -9.0% $\downarrow$ -14.0% $\downarrow$ +3.6% $\uparrow$ +34.5% $\uparrow$
Word Semantics +9.6% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0%
Wrong Candidate Generation +22.1% $\uparrow$ -6.6% $\downarrow$ +6.1% $\uparrow$ +12.4% $\uparrow$ -3.1% $\downarrow$ -1.7%
Table 14: The average downstream task performance gain or loss over the original prompt when paraphrased with a specific type from one of the six groups (columns) of tasks within a certain task family (rows) for LLaMA 3 Instruct (8B). We calculate the average across all paraphrase types in one of the six paraphrase groups and across all five tasks within one of the 24 categories. Bold indicatess the highest score per paraphrase group (column). Small changes between -3% and +3% are not colored.
Task Family Morphology Syntax Lexicon Lex.-Syn. Discourse Others Answerability Classification +1.7% 0.0% 0.0% 0.0% 0.0% 0.0% Commonsense Classification +3.6% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Coreference Resolution +1.1% +2.2% 0.0% 0.0% 0.0% 0.0% Dialogue Generation +4.7% $\uparrow$ +2.2% +4.1% $\uparrow$ +0.5% +0.6% +3.0% $\uparrow$ Fill in The Blank +4.8% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Information Extraction +2.2% -8.7% $\downarrow$ -3.3% $\downarrow$ -0.8% +0.8% +3.4% $\uparrow$ Named Entity Recognition +3.2% $\uparrow$ 0.0% -0.5% 0.0% 0.0% 0.0% Program Execution -6.0% $\downarrow$ +1.5% -4.6% $\downarrow$ +5.1% $\uparrow$ 0.0% 0.0% Question Answering +4.3% $\uparrow$ +5.6% $\uparrow$ -7.1% $\downarrow$ +10.3% $\uparrow$ +9.9% $\uparrow$ +30.4% $\uparrow$ Question Generation +12.3% $\uparrow$ +7.8% $\uparrow$ -8.1% $\downarrow$ +6.5% $\uparrow$ +10.2% $\uparrow$ -3.0% Question Rewriting -4.4% $\downarrow$ -53.7% $\downarrow$ -5.7% $\downarrow$ -13.4% $\downarrow$ -4.8% $\downarrow$ -4.6% $\downarrow$ Question Understanding +9.3% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Sentence Composition -8.6% $\downarrow$ +2.6% +2.8% -0.9% -8.7% $\downarrow$ +10.6% $\uparrow$ Sentiment Analysis +2.2% 0.0% 0.0% 0.0% 0.0% 0.0% Summarization +11.7% $\uparrow$ +9.9% $\uparrow$ +8.1% $\uparrow$ +7.7% $\uparrow$ -1.8% +3.6% $\uparrow$ Text Categorization +1.5% -6.3% $\downarrow$ -0.8% +20.0% $\uparrow$ 0.0% 0.0% Text Completion +2.7% -1.1% -4.4% $\downarrow$ -0.2% +15.2% $\uparrow$ -0.2% Text Matching +1.0% 0.0% 0.0% 0.0% 0.0% 0.0% Text Quality Evaluation +4.1% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Text to Code +1.9% -4.9% $\downarrow$ -2.0% -2.4% 0.0% -1.4% Textual Entailment +3.8% $\uparrow$ 0.0% 0.0% 0.0% 0.0% 0.0% Title Generation +7.4% $\uparrow$ +4.1% $\uparrow$ +7.3% $\uparrow$ -7.5% $\downarrow$ +5.0% $\uparrow$ 0.0% Word Semantics +2.5% 0.0% 0.0% 0.0% 0.0% 0.0% Wrong Candidate Generation +26.8% $\uparrow$ -6.6% $\downarrow$ +8.0% $\uparrow$ -4.7% $\downarrow$ +1.6% +21.4% $\uparrow$
Table 15: The average downstream task performance gain or loss over the original prompt when paraphrased with a specific type from one of the six groups (columns) of tasks within a certain task family (rows) for LLaMA 3 Instruct (70B). We calculate the average across all paraphrase types in one of the six paraphrase groups and across all five tasks within one of the 24 categories. Bold indicatess the highest score per paraphrase group (column). Small changes between -3% and +3% are not colored.
Citation for this Paper